OpenStudy (macabbey101):

HELP PLEASE!:) I GIVE MEDALS :) What value of N solves the equation? 3^N = 1/81 A. –243 B. –27 C. –4 D. 4

OpenStudy (anonymous):

plug in each value, use a calculator.

OpenStudy (anonymous):

plug in -4... what do you get @macabbey101 ?

OpenStudy (macabbey101):

its not working for me :/ its just giving me a bunch of numbers..

OpenStudy (macabbey101):

is it D?

OpenStudy (anonymous):

\[3\ ^{-4}=?\]

OpenStudy (anonymous):

\[3^{-4}=\frac{ 1 }{ 81 }\]

OpenStudy (anonymous):

so what is you're answer?

OpenStudy (anonymous):

c.

OpenStudy (anonymous):

Review your exponent laws. You have to understand what a negative exponent means. Also you have to recognize (or use trial and error) to see that 81 is 3 to the exponent 4.

OpenStudy (anonymous):

\[3^n = \frac{ 1 }{ 81 } = \frac{ 1 }{ 3^{4} } = 3^{-4} \]

OpenStudy (anonymous):

So n = -4 which is answer c.